+q +q +q 2r +Q +30 +20 +O Number 1 Number 2 Number 3 9. What is true about the electric potential energy of arrangements 1 and 2? The electric potential energy of number 1 is greater than number 2 The electric potential energy of number 1 is smaller than number 2 The electric potential energy is the same in number 1 and 2, and it is positive The electric potential energy is the same in number 1 and 2, and it is negative a. b. c. d. 10. If Q 2 x 10-6 C, q-3x 10-6 C, and R 5 cm, what is the net force on the +q charge in arrangement 3? a. b. c. d. e. 86.3 N to the left 86.3 N to the right 43.2 N to the left 43.2 N to the right 10.8 N to the left
Answer #1

9) The electric potential energy of number 1 is

U_1=k(2Q)q/r=2kQq/r

The electric potential energy of number 2 is

U_2=k(Q)q/(2r)=kQq/(2r)

a) Hence the electric potential energy of number 1 is greater than number 2

10) The net force on q is

F=k\left(\frac{3Qq}{R^2}-\frac{Qq}{R^2} \right )=\frac{2kQq}{R^2}=\frac{2\times 8.99\times 10^9\times 2\times 3\times 10^{-12}}{0.05^2}=43.2 \,\,N

The direction of the force is to the left

Hence the answer is

c) 43.2 N to the left